14. The perimeter of the right-angled triangle ABC is 80 cm. Its base is 6x and one of the side ribs x². Find x.


1
Expert's answer
2021-06-02T08:43:11-0400

If one of the side ribs, which equals x² is hypotenuse, then the other rib length is: "b^2 = (x^2)^2 - (6x)^2 = x^4 - 36x^2."

"b = \sqrt{x^4 - 36x^2}"

The perimeter of this triangle is:

"P = x^2 + 6x + \sqrt{x^4 - 36x^2} = 80,"

"\sqrt{x^4 - 36x^2} = 80 - x^2 - 6x,"

"x^4 - 36x^2 = 6,400 - 80x^2 - 480x - 80x^2 + x^4 + 6x^3 - 480x + 6x^3 + 36x^2,"

"12x^3 - 88x^2 - 960x + 6,400 = 0,"

"3x^3 - 22x^2 - 240x + 1,600 = 0,"

"(3x^3 - 30x^2) + (8x^2- 240x + 1,600) = 0,"

"3x^2(x - 10) + 8(x^2 - 30x + 200) = 0,"

"3x^2(x - 10) + 8(x - 10)(x - 20) = 0,"

"(x - 10)(3x^2 + 8x - 160) = 0,"

So, x1 = 10 is one of the roots of this equation.

D = 64 + 1920 = 1984.

"x2 = \frac{-8 + \sqrt{1984}}{6} = 6.09,"

"x3 = \frac{-8 - \sqrt{1984}}{6} = -8.76," which is not possible in our case.